Here is an example for importing graphs Using the Geom style.

∫_{0}^{1} f  μ_{2} lim_{n→∞} x_{n} = 0  here ·

b˙ new Figure 1. The title. This is the caption

The position of bounding box was shown by the four corners. If you did not specify the right coordinate of the bounding box, this facility can help you to detect this sort of error. For example

μ_{2}

bFigure 2. Place caption here.

To turn of this facility, use the command: \proofingfalse.